OpenStudy (anonymous):

Make 'c' the subject 3c+b ----- =c+a 2

OpenStudy (anonymous):

first step \[ \frac{2(3c+b)}{2}=2(c+a) \]

OpenStudy (anonymous):

3c + b = 2c + 2a 3c - 2c = 2a - b c = 2a - b

OpenStudy (anonymous):

Thanks a lot